# Survey on Publicly Available Sinhala Natural Language Processing Tools and Research

### v24

@@ 3.14 Word Sense Disambiguation
There have been multiple attempts to do word sense disambiguation (WSD) [322, 323, 324, 325, 326] for Sinhala. For this, Arukgoda et al. [191] have proposed a system named Aruth based on the Lesk Algorithm[327]. An online tool125125125http://aruth.herokuapp.com/, an API126126126https://bit.ly/3sJEYbS of the algorithm, and code along with data on github127127127https://github.com/jseanm1/aruthSWSD are available. For the same task, Marasinghe et al. [328] have proposed a system based on probabilistic modeling. A dialogue act recognition system which utilizes simple classification algorithms has been proposed by Palihakkara et al. [329]. A word sense disambiguation tool named Sinsense was introduced by Subasingha [330]. They used cross-lingual sense disambiguation where English sense disambiguation was used to obtain Sinhala sense disambiguation. However, neither their tool nor their full research is publicly available.
@@ 3.15 Text Summarizing
A deterministic process flow for automatic Sinhala text summarizing was proposed by Welgama [331]. The study by Wimalasuriya [332], which has the same name as the above work by Welgama [331], uses the graph-based TextRank algorithm for automatic Sinhala text summarizing. The use case of Sinhala Text summarization for government gazettes was explored by Jayawardane [333]. The study by Rathnayake et al. [334] compared the results of extractive and abstractive summarization on Sinhala textbooks. The study by Jahan and Wijesekara [335] compared the abstractive summarizing methods of TF-IDF and Text-Rank for Sinhala using ROUGE as the evaluation score. They concluded TF-IDF to be the superior choice. Hasan et al. [151] introduced the XL-sum128128128https://github.com/csebuetnlp/xl-sum/tree/master dataset which includes 3,414 Sinhala documents and their summaries collected from Sinhala BBC website129129129https://www.bbc.com/sinhala. The M3LS130130130https://github.com/anubhav-jangra/M3LS multi-modal dataset introduced by Verma et al. [152] contains 10,148 Sinhala documents, relevant images, and their summaries collected from Sinhala BBC website131131131https://www.bbc.com/sinhala. Further, they claim that MT5 [336] obtains the best ROUGE-1 and ROUGE-L f scores for their Sinhala data set. The first Sinhala multi-document summarizing dataset, M2DS132132132https://huggingface.co/datasets/KushanH/m2ds, was created by Hewapathirana et al. [153]. The M2DS dataset consists of 23.5k Sinhala documents in 5.5k clusters with each cluster having a golden summary. Thier code and pre-trained models are also available on github133133133https://github.com/KushanMH/m2ds.
@@ 3.17 Phonological Tools
On the case of phonological layer, a report on Sinhala phonetics and phonology was published by Wasala and Gamage [338]Wickramasinghe et al. [339] discussed the practical issues in developing Sinhala Text-to-Speech and Speech Recognition systems. The Massively Multilingual Speech (MMS)134134134https://github.com/facebookresearch/fairseq/tree/main/examples/mms data set created by Pratap et al. [145] has Sinhala data for the Spoken Language Identification (LID) task. The meta-study by Al-Fraihat et al. [340] compared the status of Sinhala speech recognition research against 17 other languages.
@@ 3.17.2 Speech-to-Text
On the converse, Nadungodage et al. [365] have done a series of work on Sinhala speech recognition with special notice given to Sinhala being a resource poor language. This project divides its focus on: continuity [366], active learning [367], and speaker adaptation [368]. A Sinhala speech recognition for voice dialing which is speaker independent was proposed by Amarasingha and Gamini [369] and on the other end, a Sinhala speech recognition methodology for interactive voice response systems, which are accessed through mobile phones was proposed by Manamperi et al. [370]. A Sinhala speech to Unicode text converter for the disaster relief domain was proposed by Prasangini and Nagahamulla [371]Priyadarshani [372] proposes a method for speaker dependant speech recognition based on their previous work on: dynamic time warping for recognizing isolated Sinhala words [373], genetic algorithms [374], and syllable segmentation method utilizing acoustic envelopes [375]. The method proposed by Gunasekara and Meegama [376] utilizes an HMM model for Sinhala speech-to-text. A Sinhala speech recognizer supporting bi-directional conversion between Unicode Sinhala and phonetic English was proposed by Punchimudiyanse and Meegama [377]. The work by Karunanayake et al. [378] transfer learns CNNs for transcribing free-form Sinhala and Tamil speech data sets for the purpose of classification. Dilshan [379] conducted a study for the specific use case of transcribing number sequences in continuous Sinhala speech.  Gamage et al. [380] explored the use of combinational acoustic models such as Deep Neural Network - Hidden Markov Model (DNN-HMM) [381] and Subspace Gaussian Mixture Model (SGMM) [381] in Sinhala speech recognition. In the follow-up work, Gamage et al. [382] extended that work with end-to-end Lattice-Free Maximum Mutual Information (e2e LF-MMI) model [383] which is claimed to be a viable solution for low resource language speech recognition by Carmantini et al. [384]. However, it was shown that the new model slightly underperforms compared to the state-of-the-art result. Later, they conducted further development on their model in a follow-up work [385].  Karunathilaka et al. [386] explore Sinhala speech recognition using deep learning models such as: pre-trained DNN, DNN, TDNN, TDNN+LSTM. The first half of the study by Arafath [387] dealt with recognizing Sinhala speech using LSTMs. Gamage et al. [360] included a Sinhala speech recognition module as one of the three modules present in their currency recognition system. Time-delay neural network architectures (including multistream CNN architecture) were used for acoustic modeling of Sinhala Automatic Speech Recognition (ASR) by Warusawithana et al. [388]. They have used the Kaldi speech recognition toolkit [389] for training the ASR models. As part of their child [sic] cognitive ability assessment model, Kahawanugoda et al. [390], proposed a Sinhala speech recognition system. The study by Azir et al. [391] attempts to identify number sequences spoken in Sinhala. TacoSi introduced by Arachchige and Weerasinghe [392] is based on Tacotron [393] and has been evaluated with 10 human evaluators to determine its text-to-speech quality. Nanayakkara [394], Nanayakkara and Weerasinghe [395] used DeepSpeech135135135https://github.com/mozilla/DeepSpeech by Mozilla for Sinhala speech recognition. Gunarathne et al. [396] used an earlier version of the CMUSphinx toolkit136136136https://cmusphinx.github.io/ to transcribe Sinhala speech to text. The later work by Akesh and Meegama [397] also used CMUSphinx toolkit on features extracted from Mel-frequency cepstral coefficients (MFCC) [398, 399] to automatically generate Sinhala subtitles from Speech. The work by Wickramaarachchi et al. [400] also used MFCC features but focused on Sinhala speech intonation recognition for the purpose of detecting speech impediments in young children. Dissanayaka et al. [401] have introduced Word Sri an application that is voice-activated and capable of Sinhala grammar checking and plagiarism checking. tha [402] have introduced a Sinhala text to Speech dataset137137137https://github.com/SiTa-SpeakerDiarization/SiTa transcribed from 60 videos with a total of 602 minutes. As the baseline, they report the lowest Diarization Error Rate (DER) of 6.0 for Powerset Cross-Entropy Diarization [403]. While the data set link points to a github repository, it does not contain the data set; it points to a project website138138138https://sita-speakerdiarization.github.io/ which then points to a download page139139139https://rtuthaya.staff.uom.lk/resources/dataset/44 where you need to send a request to the authors to obtain the data set.
@@ 3.17.5 Speech classification
The Sinhala speech classification system proposed by Buddhika et al. [246] does so without converting the speech-to-text. However, they report that this approach only works for specific domains with well-defined limited vocabularies. The work by Dinushika et al. [410] uses automatic speech recognition of Sinhala for speech command classification. Extending that, Kavmini et al. [411] presented a Sinhala speech command classification system which can be used for downstream applications. The voice assistant system created by Senarathne et al. [412] is capable of handling Sinhala voice commands. The work by Welarathna et al. [413] used CNNs to classify emotions (sad, disgust, surprise, neutral, happy, calm, fear, and angry) of Sinhala speech by Autistic children. Sundarapperuma [414] created a speech emotion detection system for Sinhala.
@@ 3.18.1 Printed Text
The XTREME-UP140140140https://github.com/google-research/xtreme-up data set created by Ruder et al. [144] contains a Sinhala data set for the OCR task. The data was obtained from book transcriptions. An attempt for Sinhala OCR system has been taken by Rajapakse et al. [418] before any other work has been done on the topic. Much later, a linear symmetry-based approach was proposed by Premaratne and Bigun [419, 420]. They then used hidden Markov model-based optimization on the recognized Sinhala script [421]. Similarly, Hewavitharana et al. [422] used hidden Markov models for off-line Sinhala character recognition. Herath et al. [123], Herath and Medagoda [124] developed a prepossessing engine based on a template matching for printed Sinhala OCR. Statistical approaches with histogram projections for Sinhala character recognition is proposed by Hewavitharana and Kodikara [423], by Ajward et al. [424], and by Madushanka et al. [425].  Karunanayaka et al. [426] also did off-line Sinhala character recognition with a use case for postal city name recognition. A separate group had attempted Sinhala OCR [427] mainly involving the nearest-neighbor method [428, 429]. A study by Ediriweera [430] uses dictionaries to correct errors in Sinhala OCR. An early attempt for Sinhala OCR by Dias et al. [431] has been extended to be online and made available to use via desktops [432] and hand-held devices [433] with the ability to recognize handwriting. A simple neural network based approach for Sinhala OCR was utilized by Rimas et al. [434]. A fuzzy-based model for identifying printed Sinhala characters was proposed by Gunarathna et al. [435]Premachandra et al. [436] proposes a simple back-propagation artificial neural network with hand crafted features for Sinhala character recognition. Another neural network with specialized feature extraction for Sinhala character recognition was proposed by Jayamaha and Naleer [437]. On the matter of neural networks and feature extraction, a feature selection process for Sinhala OCR was proposed by Kumara and Ragel [438].  Jayawickrama et al. [439] worked on Sinhala printed characters with special focus on handling diacritic vowels. However, they opted to refer to diacritic vowels as modifiers in their work. Gunawardhana and Ranathunga [440] proposed a limited approach to recognize Sinhala letters on Facebook images. A CNN-based methodology to improve printed Sinhala character OCR was proposed by Liyanage [441]. Printed Character Recognition (PCR) was used by Vasantharajan and Thayasivam [113] to create a large-scale Tamil-Sinhala-English parallel corpus. A meta-study on the effects of text genre, image resolution, and algorithmic complexity needed for Sinhala OCR from books and newspapers was conducted by Anuradha et al. [442].  Anuradha et al. [443] used Tesseract 3141141141https://tesseract-ocr.github.io/ [444] for Sinhala OCR. The later study by [445] improved the accuracy of Tesseract OCR engine on Sinhala from 53.22% to 86.16% for the data set they tested on.  Maduranga and Jayalal [446] used Artificial Neural Network (ANN) based on Universe of Discourse and Self Organization Map methods to recognize multi-style printed Sinhala characters. The study by De Silva et al. [447], even though ostensibly presented as translation research, is just invoking Google Cloud service for translation. Their novelty in Sinhala NLP lies in the facility provided to OCR the text from photos and documents. The work by Thamarasee and Surendra [448] used histogram-oriented gradient descriptor (HOG descriptor) [449] and SVMs to recognize 400+ variations of Sinhala characters.
@@ 3.18.2 Handwritten Text
Fernando et al. [454] claim to have created a database for handwriting recognition research in the Sinhala language and further claims that the data set is available at the National Science Foundation (NSF) of Sri Lanka. However, the paper provides no URLs and we were not able to find the data set on the NSF website either. The work by Karunanayaka et al. [455] is focused on noise reduction and skew correction of Sinhala handwritten words. A genetic algorithm-based approach for non-cursive Sinhala handwritten script recognition was proposed by Jayasekara and Udawatta [456]Nilaweera et al. [457] compare projection and wavelet-based techniques for recognizing handwritten Sinhala script.  Silva and Kariyawasam [458] worked on segmenting Sinhala handwritten characters with special focus on handling diacritic vowels. A comparative study of few available Sinhala handwriting recognition methods was done by Silva et al. [459]Silva et al. [460] uses contour tracing for isolated characters in handwritten Sinhala text. A Sinhala handwriting OCR system which utilizes zone-based feature extraction has been proposed by Dharmapala et al. [461]. The study by Walawage and Ranathunga [462] and its follow up study by Walawage [463] specifically focus on segmentation of overlapping and touching Sinhala handwritten characters. Silva and Jayasundere [464] focused on recognizing character modifiers in Sinhala handwriting. The similarly named studies by Mariyathas et al. [465] and Wasalthilake and Kartheeswaran [466], both utilize CNN to recognize Sinhala handwriting; as does the study by Weerasinghe [467]Ifhaam and Jayalal [468] used genetic algorithms to recognize Sinhala handwritten postal addresses for postal sorting. A segmentation-based approach that utilizes an n-gram model to recognize and validate Sinhala words written on touch screens was proposed by Mahesh and Priyankara [469]. They used a CNN classifier and were able to classify 19 different Sinhala characters. The study by Rowel et al. [470] frames their work as an E-Learning platform for hearing impaired children. However, their research does not contain any work done towards Sinhala sign language to be included in Section 3.23. What they do have is an OCR system that they claim to recognize letters and digits. Even there, we are only given an example of a recognised digit. Whether or not their system can recognize Sinhala letters is not explicitly discussed. As part of their child [sic] cognitive ability assessment model, Kahawanugoda et al. [390], proposed a Sinhala handwriting letter recognition system. The study by Withana and Rupasinghe [471] used Sinhala handwritten text classification to detect Dyslexia and Dysgraphia. The study by eka [472] has used CNN to recognize Sinhala handwritten text. Karunarathne et al. [473] used Gabor-initialized CNN (GCNN) on a dataset of 6000 handwritten Sinhala character images.

### v25

@@ 3.14 Word Sense Disambiguation
There have been multiple attempts to do word sense disambiguation (WSD) [335, 336, 337, 338, 339] for Sinhala. For this, Arukgoda et al. [192] have proposed a system named Aruth based on the Lesk Algorithm[340]. An online tool129129129http://aruth.herokuapp.com/, an API130130130https://bit.ly/3sJEYbS of the algorithm, and code along with data on github131131131https://github.com/jseanm1/aruthSWSD are available. For the same task, Marasinghe et al. [341] have proposed a system based on probabilistic modeling. A dialogue act recognition system which utilizes simple classification algorithms has been proposed by Palihakkara et al. [342]. A word sense disambiguation tool named Sinsense was introduced by Subasingha [343]. They used cross-lingual sense disambiguation where English sense disambiguation was used to obtain Sinhala sense disambiguation. However, neither their tool nor their full research is publicly available.
@@ 3.15 Text Summarizing
A deterministic process flow for automatic Sinhala text summarizing was proposed by Welgama [344]. The study by Wimalasuriya [345], which has the same name as the above work by Welgama [344], uses the graph-based TextRank algorithm for automatic Sinhala text summarizing. The use case of Sinhala Text summarization for government gazettes was explored by Jayawardane [346]. The study by Rathnayake et al. [347] compared the results of extractive and abstractive summarization on Sinhala textbooks. The study by Jahan and Wijesekara [348] compared the abstractive summarizing methods of TF-IDF and Text-Rank for Sinhala using ROUGE as the evaluation score. They concluded TF-IDF to be the superior choice. Hasan et al. [151] introduced the XL-sum132132132https://github.com/csebuetnlp/xl-sum/tree/master dataset which includes 3,414 Sinhala documents and their summaries collected from Sinhala BBC website133133133https://www.bbc.com/sinhala. The M3LS134134134https://github.com/anubhav-jangra/M3LS multi-modal dataset introduced by Verma et al. [152] contains 10,148 Sinhala documents, relevant images, and their summaries collected from Sinhala BBC website135135135https://www.bbc.com/sinhala. Further, they claim that MT5 [349] obtains the best ROUGE-1 and ROUGE-L f scores for their Sinhala data set. The first Sinhala multi-document summarizing dataset, M2DS136136136https://huggingface.co/datasets/KushanH/m2ds, was created by Hewapathirana et al. [153]. The M2DS dataset consists of 23.5k Sinhala documents in 5.5k clusters with each cluster having a golden summary. Thier code and pre-trained models are also available on github137137137https://github.com/KushanMH/m2ds.
@@ 3.17 Phonological Tools
On the case of phonological layer, a report on Sinhala phonetics and phonology was published by Wasala and Gamage [351]Wickramasinghe et al. [352] discussed the practical issues in developing Sinhala Text-to-Speech and Speech Recognition systems. The Massively Multilingual Speech (MMS)138138138https://github.com/facebookresearch/fairseq/tree/main/examples/mms data set created by Pratap et al. [145] has Sinhala data for the Spoken Language Identification (LID) task. The meta-study by Al-Fraihat et al. [353] compared the status of Sinhala speech recognition research against 17 other languages.
@@ 3.17.2 Speech-to-Text
On the converse, Nadungodage et al. [378] have done a series of work on Sinhala speech recognition with special notice given to Sinhala being a resource poor language. This project divides its focus on: continuity [379], active learning [380], and speaker adaptation [381]. A Sinhala speech recognition for voice dialing which is speaker independent was proposed by Amarasingha and Gamini [382] and on the other end, a Sinhala speech recognition methodology for interactive voice response systems, which are accessed through mobile phones was proposed by Manamperi et al. [383]. A Sinhala speech to Unicode text converter for the disaster relief domain was proposed by Prasangini and Nagahamulla [384]Priyadarshani [385] proposes a method for speaker dependant speech recognition based on their previous work on: dynamic time warping for recognizing isolated Sinhala words [386], genetic algorithms [387], and syllable segmentation method utilizing acoustic envelopes [388]. The method proposed by Gunasekara and Meegama [389] utilizes an HMM model for Sinhala speech-to-text. A Sinhala speech recognizer supporting bi-directional conversion between Unicode Sinhala and phonetic English was proposed by Punchimudiyanse and Meegama [390]. The work by Karunanayake et al. [391] transfer learns CNNs for transcribing free-form Sinhala and Tamil speech data sets for the purpose of classification. Dilshan [392] conducted a study for the specific use case of transcribing number sequences in continuous Sinhala speech.  Gamage et al. [393] explored the use of combinational acoustic models such as Deep Neural Network - Hidden Markov Model (DNN-HMM) [394] and Subspace Gaussian Mixture Model (SGMM) [394] in Sinhala speech recognition. In the follow-up work, Gamage et al. [395] extended that work with end-to-end Lattice-Free Maximum Mutual Information (e2e LF-MMI) model [396] which is claimed to be a viable solution for low resource language speech recognition by Carmantini et al. [397]. However, it was shown that the new model slightly underperforms compared to the state-of-the-art result. Later, they conducted further development on their model in a follow-up work [398].  Karunathilaka et al. [399] explore Sinhala speech recognition using deep learning models such as: pre-trained DNN, DNN, TDNN, TDNN+LSTM. The first half of the study by Arafath [400] dealt with recognizing Sinhala speech using LSTMs. Gamage et al. [373] included a Sinhala speech recognition module as one of the three modules present in their currency recognition system. Time-delay neural network architectures (including multistream CNN architecture) were used for acoustic modeling of Sinhala Automatic Speech Recognition (ASR) by Warusawithana et al. [401]. They have used the Kaldi speech recognition toolkit [402] for training the ASR models. As part of their child [sic] cognitive ability assessment model, Kahawanugoda et al. [403], proposed a Sinhala speech recognition system. The study by Azir et al. [404] attempts to identify number sequences spoken in Sinhala. TacoSi introduced by Arachchige and Weerasinghe [405] is based on Tacotron [406] and has been evaluated with 10 human evaluators to determine its text-to-speech quality. Nanayakkara [407], Nanayakkara and Weerasinghe [408] used DeepSpeech139139139https://github.com/mozilla/DeepSpeech by Mozilla for Sinhala speech recognition. Gunarathne et al. [409] used an earlier version of the CMUSphinx toolkit140140140https://cmusphinx.github.io/ to transcribe Sinhala speech to text. The later work by Akesh and Meegama [410] also used CMUSphinx toolkit on features extracted from Mel-frequency cepstral coefficients (MFCC) [411, 412] to automatically generate Sinhala subtitles from Speech. The work by Wickramaarachchi et al. [413] also used MFCC features but focused on Sinhala speech intonation recognition for the purpose of detecting speech impediments in young children. Dissanayaka et al. [414] have introduced Word Sri an application that is voice-activated and capable of Sinhala grammar checking and plagiarism checking. tha [415] have introduced a Sinhala text to Speech dataset141141141https://github.com/SiTa-SpeakerDiarization/SiTa transcribed from 60 videos with a total of 602 minutes. As the baseline, they report the lowest Diarization Error Rate (DER) of 6.0 for Powerset Cross-Entropy Diarization [416]. While the data set link points to a github repository, it does not contain the data set; it points to a project website142142142https://sita-speakerdiarization.github.io/ which then points to a download page143143143https://rtuthaya.staff.uom.lk/resources/dataset/44 where you need to send a request to the authors to obtain the data set.
@@ 3.17.5 Speech classification
Luger et al. [423] introduced a data set named unsupervised_peoples_speech144144144https://huggingface.co/datasets/MLCommons/unsupervised_peoples_speech that conatains 981 hours of Sinhala speech data. The Sinhala speech classification system proposed by Buddhika et al. [248] does so without converting the speech-to-text. However, they report that this approach only works for specific domains with well-defined limited vocabularies. The work by Dinushika et al. [424] uses automatic speech recognition of Sinhala for speech command classification. Extending that, Kavmini et al. [425] presented a Sinhala speech command classification system which can be used for downstream applications. The voice assistant system created by Senarathne et al. [426] is capable of handling Sinhala voice commands. Kathriarachchi and Pulasinghe [427] developed a system to classify and predict Sinhala speech clips of children into age groups. The work by Welarathna et al. [428] used CNNs to classify emotions (sad, disgust, surprise, neutral, happy, calm, fear, and angry) of Sinhala speech by Autistic children. Sundarapperuma [429] created a speech emotion detection system for Sinhala.
@@ 3.18.1 Printed Text
The XTREME-UP145145145https://github.com/google-research/xtreme-up data set created by Ruder et al. [144] contains a Sinhala data set for the OCR task. The data was obtained from book transcriptions. An attempt for Sinhala OCR system has been taken by Rajapakse et al. [433] before any other work has been done on the topic. Much later, a linear symmetry-based approach was proposed by Premaratne and Bigun [434, 435]. They then used hidden Markov model-based optimization on the recognized Sinhala script [436]. Similarly, Hewavitharana et al. [437] used hidden Markov models for off-line Sinhala character recognition. Herath et al. [123], Herath and Medagoda [124] developed a prepossessing engine based on a template matching for printed Sinhala OCR. Statistical approaches with histogram projections for Sinhala character recognition is proposed by Hewavitharana and Kodikara [438], by Ajward et al. [439], and by Madushanka et al. [440].  Karunanayaka et al. [441] also did off-line Sinhala character recognition with a use case for postal city name recognition. A separate group had attempted Sinhala OCR [442] mainly involving the nearest-neighbor method [443, 444]. A study by Ediriweera [445] uses dictionaries to correct errors in Sinhala OCR. An early attempt for Sinhala OCR by Dias et al. [446] has been extended to be online and made available to use via desktops [447] and hand-held devices [448] with the ability to recognize handwriting. A simple neural network based approach for Sinhala OCR was utilized by Rimas et al. [449]. A fuzzy-based model for identifying printed Sinhala characters was proposed by Gunarathna et al. [450]Premachandra et al. [451] proposes a simple back-propagation artificial neural network with hand crafted features for Sinhala character recognition. Another neural network with specialized feature extraction for Sinhala character recognition was proposed by Jayamaha and Naleer [452]. On the matter of neural networks and feature extraction, a feature selection process for Sinhala OCR was proposed by Kumara and Ragel [453].  Jayawickrama et al. [454] worked on Sinhala printed characters with special focus on handling diacritic vowels. However, they opted to refer to diacritic vowels as modifiers in their work. Gunawardhana and Ranathunga [455] proposed a limited approach to recognize Sinhala letters on Facebook images. A CNN-based methodology to improve printed Sinhala character OCR was proposed by Liyanage [456]. Printed Character Recognition (PCR) was used by Vasantharajan and Thayasivam [113] to create a large-scale Tamil-Sinhala-English parallel corpus. A meta-study on the effects of text genre, image resolution, and algorithmic complexity needed for Sinhala OCR from books and newspapers was conducted by Anuradha et al. [457].  Anuradha et al. [458] used Tesseract 3146146146https://tesseract-ocr.github.io/ [459] for Sinhala OCR. The later study by [460] improved the accuracy of Tesseract OCR engine on Sinhala from 53.22% to 86.16% for the data set they tested on.  Maduranga and Jayalal [461] used Artificial Neural Network (ANN) based on Universe of Discourse and Self Organization Map methods to recognize multi-style printed Sinhala characters. The study by De Silva et al. [462], even though ostensibly presented as translation research, is just invoking Google Cloud service for translation. Their novelty in Sinhala NLP lies in the facility provided to OCR the text from photos and documents. The work by Thamarasee and Surendra [463] used histogram-oriented gradient descriptor (HOG descriptor) [464] and SVMs to recognize 400+ variations of Sinhala characters.
@@ 3.18.2 Handwritten Text
Fernando et al. [469] claim to have created a database for handwriting recognition research in the Sinhala language and further claims that the data set is available at the National Science Foundation (NSF) of Sri Lanka. However, the paper provides no URLs, and we were not able to find the dataset on the NSF website either. Weraduwa et al. [470] claims to have created a dataset of 373 Sinhala handwriting images from 84 participants where 73 are dysgraphic and 300 are non-dysgraphic. Their objective isthat this dataset may be used to detect dysgraphia in children.
@@
The work by Karunanayaka et al. [471] is focused on noise reduction and skew correction of Sinhala handwritten words. A genetic algorithm-based approach for non-cursive Sinhala handwritten script recognition was proposed by Jayasekara and Udawatta [472]Nilaweera et al. [473] compare projection and wavelet-based techniques for recognizing handwritten Sinhala script.  Silva and Kariyawasam [474] worked on segmenting Sinhala handwritten characters with special focus on handling diacritic vowels. A comparative study of few available Sinhala handwriting recognition methods was done by Silva et al. [475]Silva et al. [476] uses contour tracing for isolated characters in handwritten Sinhala text. A Sinhala handwriting OCR system which utilizes zone-based feature extraction has been proposed by Dharmapala et al. [477]. The study by Walawage and Ranathunga [478] and its follow up study by Walawage [479] specifically focus on segmentation of overlapping and touching Sinhala handwritten characters. Silva and Jayasundere [480] focused on recognizing character modifiers in Sinhala handwriting. The similarly named studies by Mariyathas et al. [481] and Wasalthilake and Kartheeswaran [482], both utilize CNN to recognize Sinhala handwriting; as does the study by Weerasinghe [483]Ifhaam and Jayalal [484] used genetic algorithms to recognize Sinhala handwritten postal addresses for postal sorting. A segmentation-based approach that utilizes an n-gram model to recognize and validate Sinhala words written on touch screens was proposed by Mahesh and Priyankara [485]. They used a CNN classifier and were able to classify 19 different Sinhala characters. The study by Rowel et al. [486] frames their work as an E-Learning platform for hearing-impaired children. However, their research does not contain any work done towards Sinhala sign language to be included in Section 3.23. What they do have is an OCR system that they claim to recognise letters and digits. Even there, we are only given an example of a recognised digit. Whether or not their system can recognise Sinhala letters is not explicitly discussed. As part of their child [sic] cognitive ability assessment model, Kahawanugoda et al. [403], proposed a Sinhala handwriting letter recognition system. The study by Withana and Rupasinghe [487] used Sinhala handwritten text classification to detect Dyslexia and Dysgraphia. The study by eka [488] has used CNN to recognize Sinhala handwritten text. Karunarathne et al. [489] used Gabor-initialised CNN (GCNN) on a dataset of 6000 handwritten Sinhala character images.